OpenStudy (anonymous):

8to the 6th pwr divided by 8 to the second pwr u keep the 8 but do you multiply,divide,subtract,or add the exponents?

OpenStudy (anonymous):

\[{8^6 \over 8^2} = {8*8*8*8*8*8 \over 8*8} = 8^{6-2} = 8^4\]

OpenStudy (anonymous):

\[{a^x \over a^y} = a^{x-y}\]\[(a^x)(a^y) = a^{x+y}\]\[(a^x)^y = a^{xy}\]\[\sqrt[y]{a^x} = (a^{x})^{1 \over y} = a^{ x \over y } \]

OpenStudy (anonymous):

does that help ?

OpenStudy (anonymous):

choices are a.8 to the 8th pwr b.8 to the 12th pwr c.8 to the 4th power d. 8 to the 3rd power